| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  53  
 | 
Fri May  7 17:35:22 2021 | 
Kanika, Sergey, Yuri, Wolfram, Michael, Shahab | Runs | Energy change +6 MeV/u | We will now go up in energy to 447 MeV (+6 
MeV) from the original energy
 
 | 
   | 
| 
  52  
 | 
Fri May  7 17:12:52 2021 | 
Kanika, Sergey, Yuri, Wolfram, Michael, Shahab | Runs | trigger time back to 50% | changed trigger time back to 50%
 
 
last file with 1% trigger: 
 | 
   | 
| 
  51  
 | 
Fri May  7 17:12:14 2021 | 
Kanika, Sergey, Yuri, Wolfram, Michael, Shahab | Runs | Energy change -6 MeV/u | changed the energy down to 435 MeV/u | 
   | 
| 
  50  
 | 
Fri May  7 17:02:51 2021 | 
Kanika, Sergey, Yuri, Wolfram, Michael, Shahab | Runs | new sextupole setting 2 | There is new sextupole settings, but the scrapers 
remain like last night (Helmut setting last 
night.) | 
   | 
| 
  49  
 | 
Fri May  7 16:17:21 2021 | 
Sergey, Yuri, Wolfram, Shahab | Runs | new sextupole setting | new sextupole setting, isomer not visible 
yet. | 
   | 
| 
  48  
 | 
Fri May  7 16:09:50 2021 | 
Yuri | Accelerator | Stop, change of sextuple settings |   | 
   | 
| 
  47  
 | 
Fri May  7 14:33:23 2021 | 
Andrew Ragan Yuri  | General | Beam is back |   | 
   | 
| 
  46  
 | 
Fri May  7 14:06:02 2021 | 
Yuri | Accelerator | SIS kicker failure |   | 
   | 
| 
  45  
 | 
Fri May  7 12:37:35 2021 | 
Ragan, Andrew, Yuri, Alex | General | Back with beam  | We are back with the beam in the ESR since 
12:30
 
ALL SCRAPERS ARE IN THE ORIGINAL POSITIONS
 | 
   | 
| 
  44  
 | 
Fri May  7 11:40:20 2021 | 
Ragan, Andrew, Yuri | General | Scraper is fixed - optimisation stage | Scraper is fixed, we start again. Scraper 
open to let everything through | 
   | 
| 
  43  
 | 
Fri May  7 10:46:40 2021 | 
Yuri | Accelerator | GE02DS3HG | The scraper is blocked at -10 mm and -unfortunately- 
blocks the beam in the ESR. Uli Popp is taking 
care. | 
   | 
| 
  42  
 | 
Fri May  7 10:05:06 2021 | 
Yuri | Accelerator | E02DS3HG does not drive | E02DS3HG stays at -35 mm and does not react | 
   | 
| 
  41  
 | 
Fri May  7 09:58:04 2021 | 
Yuri | Accelerator | Setting up KO extraction for Caves | SIS optimisation requires us to stop. Only 
sporadic shots. | 
   | 
| 
  40  
 | 
Fri May  7 09:56:43 2021 | 
Yuri | Calibration | Identification Verification | With closed scrapers there are only a few 
weak lines.
 
Open temporarily all scrapers in the ESR | 
   | 
| 
  38  
 | 
Fri May  7 09:36:32 2021 | 
Ragan, Andrew, Yuri | General | Beam stop at 0928, adjust scope | It was requested from the SIS to stop the 
beam at 09:28 
 
Yuri is adjusting the scope in order to avoid | 
   | 
| 
  37  
 | 
Fri May  7 08:55:33 2021 | 
Ragan, Andrew, Yuri | General | Trigger changed | Trigger for both the spectrum analyzers is 
changed from 50% to 1% by Yuri.
 
For this, the ESR pattern was stopped at | 
    | 
| 
  36  
 | 
Fri May  7 08:04:22 2021 | 
Carlo Bruno, Desislava Kalaydjieva | Analysis | 7:00 - 14:00 analysis shift | (See
 
https://triumfoffice365-my.sharepoint.com/:w:/g/personal/dillmann_triumf_ca/Eb_LCp7yTx5Bgc8TO8khJmEBWeS-ycdPRnG02wQout-fgQ?rtime=3n3A3xMR2Ug
 
for remote shifter instructions. To be updated/corrected | 
 15x   | 
| 
  35  
 | 
Fri May  7 07:52:12 2021 | 
Ragan, Andrew | General | To change trigger position | Xiangcheng suggested to change the trigger 
position in order to start the file with 
the injection. Nothing is changed at the | 
   | 
| 
  34  
 | 
Fri May  7 07:02:33 2021 | 
Guy Leckenby | Analysis | Online Monitoring Plots | In compliment to Liliana's plots, this is 
the summed spectra I calculated from 0338 
to 0507. | 
    | 
| 
  33  
 | 
Fri May  7 06:53:56 2021 | 
Akash, Liliana | Analysis | Some plots | Some plots of statistics up to now | 
 6x   |